# Learning What Matters: Prioritized Concept Learning via Relative Error-driven Sample Selection

### v1

@@ 5 Conclusion
6 Limitations
While PROGRESS effectively orders and prioritizes more informative skills, it randomly samples within each selected skill cluster without ranking samples by usefulness. Additionally, the accuracy-based variant incurs extra inference time to compute skill-level progress (see Appendix A), though our loss-based variant avoids this issue. However, overall, PROGRESS outperforms prior approaches while requiring no additional reference VLM and significantly less supervised data.
Acknowledgments
We acknowledge the valuable feedback provided by Aditya Chinchure, Le Zhang, Shravan Nayak, and Rabiul Awal on the early draft. The technical support extended by the Mila IDT team in managing the computational infrastructure is greatly appreciated. Additionally, Aishwarya Agrawal and Leonid Sigal received support from the Canada CIFAR AI Chair award throughout this project.
@@ Acknowledgments
  • Additional details on the Baselines used, including setup and implementation (extending Sec.4.1 and Fig. 2 of the main manuscript).
@@
  • Further Ablation studies on the Effect of Hyperparameters and Wall-clock Time comparisons (extending Sec.4.3).
@@
Appendix A Details of Experimental Setups
A.1 Baselines
@@
We follow the standard experimental settings and implementation protocols for all baselines as established in recent prior work (see COINCIDE (Lee et al., 2024)), using official code. For completeness, clarity and reproducibility, we additionally provide detailed descriptions of each baseline here.
@@
COINCIDE (Lee et al., 2024) is a strong coreset selection method that leverages concept-based clustering and mutual transferability between clusters to guide sample selection. It performs coreset selection only once before training by clustering internal activations of a separately trained additional reference VLM (e.g., TinyLLaVA(Zhou et al., 2024)). It relies on static selection strategy that does not adapt to the model’s learning progress, requires an additional pretrained VLM, ground-truth annotations for full dataset to extract activation maps, and manual intervention to select appropriate activation layers—making the method resource-intensive and difficult to scale.
@@
CLIP-Score (Hessel et al., 2022) It ranks image-instruction pairs based on visual-textual similarity computed by the CLIP model (Radford et al., 2021), selecting top-scoring samples for training. While this approach assumes that higher similarity indicates greater informativeness, it relies on static, precomputed metrics that do not adapt to the model’s learning progress. Prior work (Lee et al., 2024) has shown that such metrics often fail to capture important data modes, resulting in reduced diversity and suboptimal generalization—limitations that PROGRESS overcomes through dynamic, progress-driven selection.
@@
EL2N (Paul et al., 2021) ranks training samples based on the expected L2-norm of prediction error: $\mathbb{E}[\|p(x)-y\|_{2}]$ where $p(x)$ is the token distribution predicted by a reference VLM, $x$ is the input, and $y$ is the ground-truth label. This score reflects how confidently and accurately the reference model predicts each sample. However, it requires access to a fully trained additional VLM and ground-truth labels for the entire dataset, making it resource-intensive and static.
@@
Perplexity (Marion et al., 2023) measures the uncertainty in the model’s predictions and is defined as $\exp(-\mathbb{E}[\log p(x)])$, where $p(x)$ denotes the likelihood assigned to input $x$ by a additional reference VLM model. Samples from the middle of the perplexity distribution are selected, following prior work (Lee et al., 2024). However, it requires access to a fully trained additional VLM and, like other static metrics, often fails to capture important data modes—potentially limiting diversity and downstream generalization.
@@
SemDeDup (Abbas et al., 2023) aims to reduce redundancy by removing semantically duplicated samples. It clusters the output embeddings of the final token from a reference model’s last layer and retains a diverse subset by eliminating near-duplicates and reducing redundancy. This method also requires an additional reference VLM to extract the final token features and ground-truth labels for the entire dataset.
@@
D2-Pruning (Maharana et al., 2024) constructs a graph over training data where nodes encode sample difficulty and edges capture pairwise similarity. It selects a diverse coreset by pruning this graph while preserving representative and challenging samples. Difficulty is measured using the AUM score, defined as $p_{y}(x)-\max_{i\neq y}p_{i}(x)$, where $p_{y}(x)$ is the model’s confidence for the ground-truth label $y$. Similarity is computed using the L2 distance between average final-layer token embeddings from a additional reference VLM. This method requires access to requires an additional reference VLM for embedding extraction and scoring and ground-truth labels for the entire dataset.
@@
Self-Sup (Sorscher et al., 2022) clusters data using averaged output embeddings from the final-layer tokens of a reference model. It assigns scores based on distance to cluster centroids, selecting samples that are closest—assumed to be the most prototypical representatives of the data distribution. This method also requires access to an additional reference VLM for embedding extraction.
@@
Self-Filter (Chen et al., 2024a) is a recent coreset selection method originally proposed for the LLaVA-158k dataset (containing three vision-language tasks). It jointly fine-tunes a scoring network alongside the target VLM on the entire labeled dataset, using this as learned reference model to score and filter training samples—hence it requires an additional reference model trained on full data with full annotations. Following previous work (Lee et al., 2024), we adopt the stronger variant that also incorporates CLIP scores and features.
@@
Random. We additionally report results for Random, which finetunes the model using a coreset selected via random sampling. Despite its simplicity, Random serves as a strong and competitive baseline—prior work (Lee et al., 2024)has shown that random sampling often preserves sample diversity and can outperform more complex selection methods in certain settings.
@@
Note: We use standard setup for the baseline implementations as described prior work (see COINCIDE Appendix). For COINCIDE, EL2N, SemDeDup, D2-Pruning, and Self-Sup, we use image, question, and ground-truth answer for full dataset as inputs along with additional reference VLM (i.e TinyLLaVA) to extract features following prior work (Lee et al., 2024). Self-Filter requires full dataset to finetune additional reference network—the score-net. As a result, these baselines require an additional reference vision-language model or full dataset annotations (100%) or both.
A.2 Implementation Details.
@@
Method LLaVA-1.5 Vision-Flan
CLIP-Score high score selected high score selected
EL2N medium score selected medium score selected
Perplexity medium score selected medium score selected
SemDeDup $\bm{K}$ : 10,000 $\bm{K}$ : 5,000
D2-Pruning $\bm{k}$ : 5, $\gamma_{r}$ : 0.4, $\gamma_{f}$ : 1.0 $\bm{k}$ : 5, $\gamma_{r}$ : 0.4, $\gamma_{f}$ : 1.0
Self-Sup $\bm{K}$ : 10,000 $\bm{K}$ : 5,000
Self-Filter $\bm{k}$ : 10, $\gamma$ : 1 $\bm{k}$ : 10, $\gamma$ : 1
COINCIDE $\bm{K}$ : 10,000, $\tau$ : 0.1 $\bm{K}$ : 5,000, $\tau$ : 0.1
PROGRESS
Warmup Stage
Number of Clusters $\bm{K}$ : 10,000 $\bm{K}$ : 5,000
Warmup Ratio (w.r.t full data) 9% 8.4%
Prioritized Concept Learning
Number of Clusters $\bm{K}$ : 1,000 $\bm{K}$ : 200
Temperature of Softmax $\tau$ : 1.0 $\tau$  : 1.0
BatchSize 128 128
Selection Gap $\bm{\gamma}*BatchSize$ : 7,500 $\bm{\gamma}*BatchSize$ : 3,500
Random Exploration $\delta\%$ :  10 % $\delta\%$ :  10 %
@@
We first partition the unlabeled data pool $\mathbb{U}$ into $K$ skill clusters using spherical k-means, following the fully unsupervised concept categorization procedure described in Section 3.1. Training begins with a brief warmup phase (see details in Appendix A.3), which equips the model with basic instruction-following capability and ensures that skill-level performance estimates are reliable in the beginning of training.
@@
Hyperparameters for Baselines and PROGRESS
@@
To ensure fair comparison, we use the same hyperparameters as COINCIDE (Lee et al., 2024) for all baselines. The hyperparameters for both the baselines and PROGRESS are summarized in Table 4. For model training, we apply LoRA (Hu et al., 2021) to LLaVA-v1.5 and follow the official fine-tuning settings provided in the LLaVA-1.5 release. For Qwen2-VL, we perform full fine-tuning using the official hyperparameters specified by Qwen2-VL. For accuracy estimation, we use LLMs such as InternLM2-Chat-20B (Cai & et al., 2024) as the judge. We provide the question, ground-truth answer, and predicted answer (without the image) as input and ask the LLM to decide whether the prediction is correct. The full prompt is shown below in Table A.2. Ablation studies on all hyperparameters are provided in Section 4.3 and Appendix B.
@@
Prompt for Accuracy Estimation Given an input question and two answers: a candidate answer and a reference answer, determine if the candidate answer is correct or incorrect. Rules: The candidate answer is correct if it is semantically equivalent to the reference answer, even if they are phrased differently. The candidate answer should be marked as incorrect if it: Contains factual errors compared to the reference answer Only partially answers the question Includes hedging language (e.g., "probably", "likely", "I think", etc.) Answers a different question than what was asked Give a reason for your prediction. Output Format: Answer - correct or incorrect Reason -
A.3 Warmup Phase
@@
Following prior work (Xia et al., 2024; Wu et al., 2025), we begin with a brief warmup phase using a small subset—9% of the total dataset pool size (see ablation in Fig. 9 (a))—to equip the model with basic instruction-following capability and ensure that skill-level performance estimates are credible at the start of training—when the model is still untrained. These early estimates are crucial for tracking relative learning progress across skills in subsequent training phases.
To be effective, the warmup set should provide broad skill coverage across diverse clusters and include transferable examples that support generalization to unseen skills that model will later learn. To this end, we adopt the sampling strategy from Lee et al. (2024), selecting samples from each cluster proportionally to: $P_{i}\propto\exp\left({S_{i}}/{\tau D_{i}}\right)$ where $S_{i}$ denotes the cluster’s transferability and $D_{i}$ its density. This prioritizes clusters that are both diverse and likely to generalize well, ensuring a representative warmup set.
Importantly, the entire warmup set selection used for our method utilizes clusters generated using concatenated DINO and BERT derived features (as explained in Sec. 3.1)—instead of requiring additional reference VLMs features or ground-truth answers. As shown in Appendix B, we validate that training solely on this warmup set (without our Prioritized concept learning module in Section 3.2) yields significantly worse performance compared to our proposed strategy—highlighting the importance of our progress-driven sampling.
A.4 Comparison with ICONS
ICONS (Wu et al., 2025) is a concurrent unpublished work that differs significantly from our approach. It requires (1) high memory and compute resources—reportedly over 100 GPU hours—to compute and store gradient-based influence scores for selection, and (2) access to explicit knowledge of the target task or its distribution in the form of labeled samples from validation set of target benchmarks. This assumption is impractical in general-purpose VLM training, where target tasks may be unknown at training time and usage of such high-compute refutes the goal of efficient learning. As such, ICONS is not directly comparable and falls outside the scope of our setting, which avoids both gradient-based selection and downstream task knowledge from target benchmarks prior to training the VLM model.
@@ A.4 Comparison with ICONS
Method VQAv2 GQA VizWiz SQA-I TextVQA POPE MME MMBench LLaVA- SEED AI2D ChartQA CMMMU Rel. (%)
@@
3   Loss as Obj. 75.7 58.6 49.6 70.1 55.1 86.3 1498.4 62.5 55.5 65.5 63.4 53.3 17.3 23.7 98.4
4    Accuracy as Obj. 75.2 58.8 53.4 69.9 55.1 85.9 1483.2 61.1 54.4 65.5 63.0 52.8 17.3 24.6 98.8
Appendix B Further Ablation Studies and Analysis
B.1 Ablation Studies

### v2

@@ 5 Conclusion
  • Additional details on the Baselines (extending Sec.4.1 and Fig. 1 of the main manuscript).
@@
  • Further Ablation studies for Hyperparameters (extending Sec.4.3)
  • Analysis on Why PROGRESS Works: Representativeness of Selected Subsets and Order of Skills
@@
  • Intuition and Justification for using relative improvement (extending Sec.3).
  • In-depth analysis Wall-clock Time, Time Breakdown Analysis, Annotation Time Analysis, and Scaling Performance (extending Sec.4.3).
@@
  • Limitations of our method and LLM usage disclosure statement
6 Details of Experimental Setups
6.1 Baselines
@@
We follow the standard experimental settings and implementation protocols for all baselines as established in recent prior work (see COINCIDE [19]), using official code. For completeness, clarity and reproducibility, we additionally provide detailed descriptions of each baseline here.
@@
COINCIDE [19] is a strong coreset selection method that leverages concept-based clustering and mutual transferability between clusters to guide sample selection. COINCIDE performs coreset selection only once, before training, by clustering internal activations of a separately trained (additional) reference VLM (e.g., TinyLLaVA[50]). COINCIDE relies on static selection strategy that does not adapt to the model’s learning progress, requires an additional pretrained VLM, ground-truth annotations for full dataset to extract activation maps, and manual intervention to select appropriate activation layers—making the method resource-intensive and difficult to scale.
@@
CLIP-Score [13] It ranks image-instruction pairs based on visual-textual similarity computed by the CLIP model [37], selecting top-scoring samples for training. While this approach assumes that higher similarity indicates greater informativeness, it relies on static, precomputed metrics that do not adapt to the model’s learning progress. Prior work [19] has shown that such metrics often fail to capture important data modes, resulting in reduced diversity and suboptimal generalization—limitations that PROGRESS overcomes through dynamic, progress-driven selection.
@@
EL2N [36] ranks training samples based on the expected L2-norm of prediction error: $\mathbb{E}[\|p(x)-y\|_{2}]$ where $p(x)$ is the token distribution predicted by a reference VLM, $x$ is the input, and $y$ is the ground-truth label. This score reflects how confidently and accurately the reference model predicts each sample. However, it requires access to a fully trained additional VLM and ground-truth labels for the entire dataset, making it resource-intensive and static.
@@
Perplexity [31] measures the uncertainty in the model’s predictions and is defined as $\exp(-\mathbb{E}[\log p(x)])$, where $p(x)$ denotes the likelihood assigned to input $x$ by a additional reference VLM model. Samples from the middle of the perplexity distribution are selected, following prior work [19]. However, it requires access to a fully trained additional VLM and, like other static metrics, often fails to capture important data modes—potentially limiting diversity and downstream generalization.
@@
SemDeDup [1] aims to reduce redundancy by removing semantically duplicated samples. It clusters the output embeddings of the final token from a reference model’s last layer and retains a diverse subset by eliminating near-duplicates and reducing redundancy. This method also requires an additional reference VLM to extract the final token features and ground-truth labels for the entire dataset.
@@
D2-Pruning [29] constructs a graph over training data where nodes encode sample difficulty and edges capture pairwise similarity. It selects a diverse coreset by pruning this graph while preserving representative and challenging samples. Difficulty is measured using the AUM score, defined as $p_{y}(x)-\max_{i\neq y}p_{i}(x)$, where $p_{y}(x)$ is the model’s confidence for the ground-truth label $y$. Similarity is computed using the L2 distance between average final-layer token embeddings from a additional reference VLM. This method requires access to an additional reference VLM for embedding extraction and scoring and ground-truth labels for the entire dataset.
@@
Self-Sup [40] clusters data using averaged output embeddings from the final-layer tokens of a reference model. It assigns scores based on distance to cluster centroids, selecting samples that are closest—assumed to be the most prototypical representatives of the data distribution. This method also requires access to an additional reference VLM for embedding extraction.
@@
Self-Filter [6] is a recent coreset selection method originally proposed for the LLaVA-158k dataset (containing three vision-language tasks). It jointly fine-tunes a scoring network alongside the target VLM on the entire labeled dataset, using this as learned reference model to score and filter training samples—hence it requires an additional reference model trained on full data with full annotations. Following previous work [19], we adopt the stronger variant that also incorporates CLIP scores and features.
@@
Random. We additionally report results for Random, which finetunes the model using a coreset selected via random sampling. Despite its simplicity, Random serves as a strong and competitive baseline—prior work [19]has shown that random sampling often preserves sample diversity and can outperform more complex selection methods in certain settings.
@@
Note: We use standard setup for the baseline implementations as described in prior work (see COINCIDE Appendix [19]). For COINCIDE, EL2N, SemDeDup, D2-Pruning, and Self-Sup, we use image, question, and ground-truth answer for full dataset as inputs along with additional reference VLM (i.e., TinyLLaVA) to extract features following prior work [19]. Self-Filter requires full dataset to finetune additional reference network—the score-net. As a result, these baselines require an additional reference vision-language model or full dataset annotations (100%) or both.
Active Learning Nature of the Compared Baselines. AL selects which unlabeled samples to label next using acquisition function a(x) [44]. Higher a(x) (e.g., uncertainty, error, diversity, representativeness) implies higher expected utility of sample. EL2N [36] and Perplexity [31] are uncertainty-based AL methods with error/uncertanity as acquisition functions. Self-Filter [6] uses deep network as acquisition function. It jointly trains a deep score network with the target model to later rank and select samples. COINCIDE [19] is hybrid sampling based AL which uses a hybrid of skill transferability + diversity for acquisition of samples to select a representative coreset. Self-Sup [40] aligns with representativeness-based AL coreset selection, aiming to pick prototypical samples that best cover the data distribution.
Curriculum Learning Nature of Compared Baselines. Curriculum Learning (CL) methods can be partitioned into two types: (Type 1) External model decides difficulty and select samples; (Type 2) Models own feedback decides difficulty and select samples.
Self-Filter [6] is an instantiation of Type 1 CL methods where an external model (i.e deep score network) is jointly trained with the target model to quantify the learning difficulty of sample and select hard-examples for training. The auxiliary score network guides selection of samples expected to have significant value for model training based on difficulty.
We also compare with Type 2 curriculum policies in Table 3: 1) Easiest Selection (Table 3 row 3) - model self-evaluates and chooses easy-samples with highest absolute performance, 2) Medium difficulty Selection (Table 3, row 4)- model self-evaluates and chooses samples with medium absolute performance, 3) Hardest Selection (Table 3, row 5) - model self-evaluates and chooses hard-samples with lowest absolute performance 4) Relative-Improvement based selection (ours, Table 3, row 6-7).
Overall, our PROGRESS method (relative improvement based policy) performs much better than other CL baselines.
6.2 Implementation Details
@@
Method LLaVA-1.5-665K Vision-Flan CLIP-Score high score selected high score selected EL2N medium score selected medium score selected Perplexity medium score selected medium score selected SemDeDup $\bm{K}$ : 10,000 $\bm{K}$ : 5,000 D2-Pruning $\bm{k}$ : 5, $\gamma_{r}$ : 0.4, $\gamma_{f}$ : 1.0 $\bm{k}$ : 5, $\gamma_{r}$ : 0.4, $\gamma_{f}$ : 1.0 Self-Sup $\bm{K}$ : 10,000 $\bm{K}$ : 5,000 Self-Filter $\bm{k}$ : 10, $\gamma$ : 1 $\bm{k}$ : 10, $\gamma$ : 1 COINCIDE $\bm{K}$ : 10,000, $\tau$ : 0.1 $\bm{K}$ : 5,000, $\tau$ : 0.1 PROGRESS Warmup Stage Number of Clusters $\bm{K}$ : 10,000 $\bm{K}$ : 5,000 Warmup Ratio (w.r.t full data) 9% 8.4% Prioritized Concept Learning Number of Clusters $\bm{K}$ : 1,000 $\bm{K}$ : 200 Temperature of Softmax $\tau$ : 1.0 $\tau$  : 1.0 BatchSize 128 128 Selection Gap $\bm{\gamma}*BatchSize$ : 7,500 $\bm{\gamma}*BatchSize$ : 3,500 Random Exploration $\delta\%$ :  10 % $\delta\%$ :  10 %
@@
We first partition the unlabeled data pool $\mathbb{U}$ into $K$ skill clusters using spherical k-means, following the fully unsupervised concept categorization procedure described in Section 3.1. Training begins with a brief warmup phase (see details in Appendix 6.3), which equips the model with basic instruction-following capability and ensures that skill-level performance estimates are reliable in the beginning of training.
@@
Hyperparameters for Baselines and PROGRESS .
@@
To ensure fair comparison, we use the same hyperparameters as COINCIDE [19] for all baselines. The hyperparameters for both the baselines and PROGRESS are summarized in Table 5. For model training, we apply LoRA [14] to LLaVA-v1.5 and follow the official fine-tuning settings provided in the LLaVA-1.5 release. For Qwen2-VL, we perform full fine-tuning using the official hyperparameters specified by Qwen2-VL. For Qwen2.5-32B-Instruct, we apply LoRA and follow the official fine-tuning settings provided in LLaMA-Factory [49]. For accuracy estimation, we use LLMs such as InternLM2-Chat-20B [4] as the judge. We provide the question, ground-truth answer, and predicted answer (without the image) as input and ask the LLM to decide whether the prediction is correct. The full prompt is shown below. Ablation studies on all hyperparameters are provided in Section 4.3 and Appendix 7.
@@
Prompt for Accuracy Estimation Given an input question and two answers: a candidate answer and a reference answer, determine if the candidate answer is correct or incorrect. Rules: The candidate answer is correct if it is semantically equivalent to the reference answer, even if they are phrased differently. The candidate answer should be marked as incorrect if it: Contains factual errors compared to the reference answer Only partially answers the question Includes hedging language (e.g., ”probably”, ”likely”, ”I think”, etc.) Answers a different question than what was asked Give a reason for your prediction. Output Format: Answer - correct or incorrect Reason -
6.3 Warmup Phase
Warmup strategy details.
@@ 6.3 Warmup Phase
Following prior work [46, 45], we begin with a brief warmup phase using a small subset—9% of the total pool size (we show in Fig. 10(a) that PROGRESS is robust to even smaller warmup ratios, remaining close to full-data performance even when the warmup percentage is further reduced.) — to equip the model with basic instruction-following capability and to obtain credible skill-level performance estimates at the start of training, when the model is still untrained. These initial estimates are essential for tracking relative learning progress across skills in subsequent phases. To make the warmup set effective, it should broadly cover diverse skill clusters.
To this end, we use a simple cluster-based sampler that prioritizes clusters which are both diverse and likely to generalize well. Concretely, we sample from each unsupervised concept cluster with probability $P_{i}\propto\exp\left({S_{i}}/{\tau D_{i}}\right)$ where $S_{i}$ and $D_{i}$ denote the cluster’s transferability and density, respectively, and $\tau$ is temperature following [19]. Importantly, unlike  [19], our clusters and scores are computed directly from concatenated self-supervised DINO and BERT features (Sec. 3.1), and do not require any additional auxiliary reference VLM (i.e TinyLLaVA) features or ground-truth answers.
Table 4 (main manuscript) shows that the warmup phase alone provides only modest performance (94.6% Rel.), while PROGRESS yields large gains over this baseline (+4% relative, rows 2 vs. 1). This demonstrates that warmup merely initializes reasonable per-cluster estimates, whereas the major improvements stem from PROGRESS ’s progress-driven sample selection. Moreover, row 3 shows that PROGRESS remains robust to different warmup strategies, achieving near–full-data performance even with a weak (random sampling) warmup.
@@
[…]
@@
[…]
6.4 Comparison with ICONS
ICONS [45] is a concurrent unpublished work that differs significantly from our approach. It requires (1) high memory and compute resources—reportedly over 100 GPU hours—to compute and store gradient-based influence scores for selection, and (2) access to explicit knowledge of the target task or its distribution in the form of labeled samples from validation set of target benchmarks. This assumption is impractical in general-purpose VLM training, where target tasks may be unknown at training time and usage of such high-compute refutes the goal of efficient learning. As such, ICONS is not directly comparable and falls outside the scope of our setting, which avoids both gradient-based selection and downstream task knowledge from target benchmarks prior to training the VLM model.
6.5 Intuition, justification, and grounding for using relative improvement and softmax-based sampling
Empirical Validation. We show ablation results and discussion in Table 3 for different selection policy including Our Relative improvement strategy (Table 3, row 6,7) vs Easy, Medium and Hard selection based on hard-thresholding on absolute performance. Overall, our relative improvement policy performs much better than other variants
Intuition, justification for using relative improvement. Inspired by prior curriculum literature [34, 38, 3] we use relative improvement over absolute gains because it scales progress by a skill’s baseline performance i.e normalizes across tasks of different difficulty and scale: tasks that are too easy (high baseline) or too hard (low baseline) tend to have small relative improvements, whereas tasks at a “moderate” difficulty yield the largest relative improvements and therefore get prioritized. This aligns with the zone‑of‑proximal‑development [42, 17], problems that are neither too easy nor too hard produce the largest advantage, focusing on learning where progress is most productive.
In contrast, using absolute improvement would bias selection towards tasks with high raw score over‑favoring raw score jumps and fail to account for the fact that a one‑point increase on a hard skill is more meaningful than the same gain on an easy one, fixed thresholds likewise cannot adapt to varying skill difficulties. Overall, relative improvement avoids over‑favoring weak skills with noisy raw swings and instead emphasizes areas where progress per unit of remaining error is highest, providing a stable, self‑paced curriculum [34, 38].
Intuition for Softmax Sampling. We map $\Delta_{k}^{t}$ to sampling weights using a temperature-controlled softmax, $p_{k}\propto\exp(\Delta_{k}^{t}/\tau)$, which provides a smooth, entropy-regularized trade-off between exploiting high-utility skills and exploring others. A small floor probability safeguards rare skills. This results in a stable, noise-robust scheduler that outperforms absolute gains and hard thresholds. The temperature parameter $\tau$ explicitly balances informativeness and diversity which is crucial for effective learning and avoid mode collapse (see Fig 5).
7 Further Ablation Studies and Analysis
7.1 Scalability and Generalization to Larger VLMs
Here we additionally present results for Qwen2.5-VL-32B-Instruct in Table 8, which we instruction-tuned using our method on the LLaVA-665K dataset, using a 20% sampling ratio following standard protocol. PROGRESS (trained with just 20% data) performs better than full- data finetuning, achieving 100.2% relative performance compared to full-data fine-tuning, demonstrating our method’s scalability potential and its generalization to new large scale architectures.
Table 8: Scalability and Fine-tuning. We report results for Qwen2.5-VL-32B on the LLaVA-665K dataset using 20% sampling ratio.
[…]
en cn Wild
[…]
7.2 Ablation Studies